A ray of light is incident through glass, with refractive index 1.52, on an interference separating glass and water with refractive index 1.32. What is the angle of refraction if the angle of incidence of the ray in glass is 25°?

1
Expert's answer
2022-04-11T12:06:50-0400

The Snell's law says

"n_1\sin\theta_1=n_2\sin\theta_2"

So, the angle of refraction

"\theta_2=\sin^{-1}\frac{n_1\sin\theta_1}{n_2}"

"\theta_2=\sin^{-1}\frac{1.32\sin25^\circ}{1.52}=21.5^\circ"